\providecommand{\href}[2]{#2}\begingroup\raggedright\begin{thebibliography}{10}

\bibitem{Ooguri:2018wrx}
H.~Ooguri, E.~Palti, G.~Shiu and C.~Vafa, \emph{{Distance and de Sitter
  Conjectures on the Swampland}},
  \href{https://arxiv.org/abs/1810.05506}{{\ttfamily 1810.05506}}.

\bibitem{Ooguri:2006in}
H.~Ooguri and C.~Vafa, \emph{{On the Geometry of the String Landscape and the
  Swampland}},
  \href{https://doi.org/10.1016/j.nuclphysb.2006.10.033}{\emph{Nucl. Phys.}
  {\bfseries B766} (2007) 21}
  [\href{https://arxiv.org/abs/hep-th/0605264}{{\ttfamily hep-th/0605264}}].

\bibitem{Bousso:1999xy}
R.~Bousso, \emph{{A Covariant entropy conjecture}},
  \href{https://doi.org/10.1088/1126-6708/1999/07/004}{\emph{JHEP} {\bfseries
  07} (1999) 004} [\href{https://arxiv.org/abs/hep-th/9905177}{{\ttfamily
  hep-th/9905177}}].

\bibitem{Obied:2018sgi}
G.~Obied, H.~Ooguri, L.~Spodyneiko and C.~Vafa, \emph{{De Sitter Space and the
  Swampland}},  \href{https://arxiv.org/abs/1806.08362}{{\ttfamily
  1806.08362}}.

\bibitem{Brennan:2017rbf}
T.~D. Brennan, F.~Carta and C.~Vafa, \emph{{The String Landscape, the
  Swampland, and the Missing Corner}},
  \href{https://arxiv.org/abs/1711.00864}{{\ttfamily 1711.00864}}.

\bibitem{Danielsson:2018ztv}
U.~H. Danielsson and T.~Van~Riet, \emph{{What if string theory has no de Sitter
  vacua?}},  \href{https://arxiv.org/abs/1804.01120}{{\ttfamily 1804.01120}}.

\bibitem{Agrawal:2018own}
P.~Agrawal, G.~Obied, P.~J. Steinhardt and C.~Vafa, \emph{{On the Cosmological
  Implications of the String Swampland}},
  \href{https://doi.org/10.1016/j.physletb.2018.07.040}{\emph{Phys. Lett.}
  {\bfseries B784} (2018) 271}
  [\href{https://arxiv.org/abs/1806.09718}{{\ttfamily 1806.09718}}].

\bibitem{Andriot:2018wzk}
D.~Andriot, \emph{{On the de Sitter swampland criterion}},
  \href{https://doi.org/10.1016/j.physletb.2018.09.022}{\emph{Phys. Lett.}
  {\bfseries B785} (2018) 570}
  [\href{https://arxiv.org/abs/1806.10999}{{\ttfamily 1806.10999}}].

\bibitem{Dvali:2018fqu}
G.~Dvali and C.~Gomez, \emph{{On Exclusion of Positive Cosmological Constant}},
   \href{https://arxiv.org/abs/1806.10877}{{\ttfamily 1806.10877}}.

\bibitem{Achucarro:2018vey}
A.~Achúcarro and G.~A. Palma, \emph{{The string swampland constraints require
  multi-field inflation}},  \href{https://arxiv.org/abs/1807.04390}{{\ttfamily
  1807.04390}}.

\bibitem{Garg:2018reu}
S.~K. Garg and C.~Krishnan, \emph{{Bounds on Slow Roll and the de Sitter
  Swampland}},  \href{https://arxiv.org/abs/1807.05193}{{\ttfamily
  1807.05193}}.

\bibitem{Lehners:2018vgi}
J.-L. Lehners, \emph{{Small-Field and Scale-Free: Inflation and Ekpyrosis at
  their Extremes}},  \href{https://arxiv.org/abs/1807.05240}{{\ttfamily
  1807.05240}}.

\bibitem{Kehagias:2018uem}
A.~Kehagias and A.~Riotto, \emph{{A note on Inflation and the Swampland}},
  \href{https://arxiv.org/abs/1807.05445}{{\ttfamily 1807.05445}}.

\bibitem{Dias:2018ngv}
M.~Dias, J.~Frazer, A.~Retolaza and A.~Westphal, \emph{{Primordial
  Gravitational Waves and the Swampland}},
  \href{https://arxiv.org/abs/1807.06579}{{\ttfamily 1807.06579}}.

\bibitem{Denef:2018etk}
F.~Denef, A.~Hebecker and T.~Wrase, \emph{{de Sitter swampland conjecture and
  the Higgs potential}},
  \href{https://doi.org/10.1103/PhysRevD.98.086004}{\emph{Phys. Rev.}
  {\bfseries D98} (2018) 086004}
  [\href{https://arxiv.org/abs/1807.06581}{{\ttfamily 1807.06581}}].

\bibitem{Colgain:2018wgk}
E.~O. Colgain, M.~H. P.~M. Van~Putten and H.~Yavartanoo, \emph{{$H_0$ tension
  and the de Sitter Swampland}},
  \href{https://arxiv.org/abs/1807.07451}{{\ttfamily 1807.07451}}.

\bibitem{Paban:2018ole}
S.~Paban and R.~Rosati, \emph{{Inflation in Multi-field Modified DBM
  Potentials}},
  \href{https://doi.org/10.1088/1475-7516/2018/09/042}{\emph{JCAP} {\bfseries
  1809} (2018) 042} [\href{https://arxiv.org/abs/1807.07654}{{\ttfamily
  1807.07654}}].

\bibitem{Roupec:2018mbn}
C.~Roupec and T.~Wrase, \emph{{de Sitter extrema and the swampland}},
  \href{https://arxiv.org/abs/1807.09538}{{\ttfamily 1807.09538}}.

\bibitem{Andriot:2018ept}
D.~Andriot, \emph{{New constraints on classical de Sitter: flirting with the
  swampland}},  \href{https://arxiv.org/abs/1807.09698}{{\ttfamily
  1807.09698}}.

\bibitem{Matsui:2018bsy}
H.~Matsui and F.~Takahashi, \emph{{Eternal Inflation and Swampland
  Conjectures}},  \href{https://arxiv.org/abs/1807.11938}{{\ttfamily
  1807.11938}}.

\bibitem{Ben-Dayan:2018mhe}
I.~Ben-Dayan, \emph{{Draining the Swampland}},
  \href{https://arxiv.org/abs/1808.01615}{{\ttfamily 1808.01615}}.

\bibitem{Heisenberg:2018yae}
L.~Heisenberg, M.~Bartelmann, R.~Brandenberger and A.~Refregier, \emph{{Dark
  Energy in the Swampland}},
  \href{https://arxiv.org/abs/1808.02877}{{\ttfamily 1808.02877}}.

\bibitem{Damian:2018tlf}
C.~Damian and O.~Loaiza-Brito, \emph{{Two-field axion inflation and the
  swampland constraint in the flux-scaling scenario}},
  \href{https://arxiv.org/abs/1808.03397}{{\ttfamily 1808.03397}}.

\bibitem{Conlon:2018eyr}
J.~P. Conlon, \emph{{The de Sitter swampland conjecture and supersymmetric AdS
  vacua}},  \href{https://arxiv.org/abs/1808.05040}{{\ttfamily 1808.05040}}.

\bibitem{Kinney:2018nny}
W.~H. Kinney, S.~Vagnozzi and L.~Visinelli, \emph{{The Zoo Plot Meets the
  Swampland: Mutual (In)Consistency of Single-Field Inflation, String
  Conjectures, and Cosmological Data}},
  \href{https://arxiv.org/abs/1808.06424}{{\ttfamily 1808.06424}}.

\bibitem{Dasgupta:2018rtp}
K.~Dasgupta, M.~Emelin, E.~McDonough and R.~Tatar, \emph{{Quantum Corrections
  and the de Sitter Swampland Conjecture}},
  \href{https://arxiv.org/abs/1808.07498}{{\ttfamily 1808.07498}}.

\bibitem{Cicoli:2018kdo}
M.~Cicoli, S.~De~Alwis, A.~Maharana, F.~Muia and F.~Quevedo, \emph{{De Sitter
  vs Quintessence in String Theory}},
  \href{https://arxiv.org/abs/1808.08967}{{\ttfamily 1808.08967}}.

\bibitem{Kachru:2018aqn}
S.~Kachru and S.~P. Trivedi, \emph{{A comment on effective field theories of
  flux vacua}},  \href{https://arxiv.org/abs/1808.08971}{{\ttfamily
  1808.08971}}.

\bibitem{Akrami:2018ylq}
Y.~Akrami, R.~Kallosh, A.~Linde and V.~Vardanyan, \emph{{The landscape, the
  swampland and the era of precision cosmology}},
  \href{https://arxiv.org/abs/1808.09440}{{\ttfamily 1808.09440}}.

\bibitem{Murayama:2018lie}
H.~Murayama, M.~Yamazaki and T.~T. Yanagida, \emph{{Do We Live in the
  Swampland?}},  \href{https://arxiv.org/abs/1809.00478}{{\ttfamily
  1809.00478}}.

\bibitem{Marsh:2018kub}
M.~C.~D. Marsh, \emph{{The Swampland, Quintessence and the Vacuum Energy}},
  \href{https://arxiv.org/abs/1809.00726}{{\ttfamily 1809.00726}}.

\bibitem{Choi:2018rze}
K.~Choi, D.~Chway and C.~S. Shin, \emph{{The dS swampland conjecture with the
  electroweak symmetry and QCD chiral symmetry breaking}},
  \href{https://arxiv.org/abs/1809.01475}{{\ttfamily 1809.01475}}.

\bibitem{DAmico:2018mnx}
G.~D'Amico, N.~Kaloper and A.~Lawrence, \emph{{Strongly Coupled Quintessence}},
   \href{https://arxiv.org/abs/1809.05109}{{\ttfamily 1809.05109}}.

\bibitem{Han:2018yrk}
C.~Han, S.~Pi and M.~Sasaki, \emph{{Quintessence Saves Higgs Instability}},
  \href{https://arxiv.org/abs/1809.05507}{{\ttfamily 1809.05507}}.

\bibitem{Moritz:2018ani}
J.~Moritz, A.~Retolaza and A.~Westphal, \emph{{On uplifts by warped
  anti-D3-branes}},  \href{https://arxiv.org/abs/1809.06618}{{\ttfamily
  1809.06618}}.

\bibitem{Halverson:2018cio}
J.~Halverson and F.~Ruehle, \emph{{Computational Complexity of Vacua and
  Near-Vacua in Field and String Theory}},
  \href{https://arxiv.org/abs/1809.08279}{{\ttfamily 1809.08279}}.

\bibitem{Anguelova:2018vyr}
L.~Anguelova, E.~M. Babalic and C.~I. Lazaroiu, \emph{{Two-field Cosmological
  $\alpha$-attractors with Noether Symmetry}},
  \href{https://arxiv.org/abs/1809.10563}{{\ttfamily 1809.10563}}.

\bibitem{Lin:2018kjm}
C.-M. Lin, K.-W. Ng and K.~Cheung, \emph{{Chaotic inflation on the brane and
  the Swampland Criteria}},  \href{https://arxiv.org/abs/1810.01644}{{\ttfamily
  1810.01644}}.

\bibitem{Hamaguchi:2018vtv}
K.~Hamaguchi, M.~Ibe and T.~Moroi, \emph{{The swampland conjecture and the
  Higgs expectation value}},
  \href{https://arxiv.org/abs/1810.02095}{{\ttfamily 1810.02095}}.

\bibitem{Ashoorioon:2018sqb}
A.~Ashoorioon, \emph{{Rescuing Single Field Inflation from the Swampland}},
  \href{https://arxiv.org/abs/1810.04001}{{\ttfamily 1810.04001}}.

\bibitem{Wang:2018kly}
S.-J. Wang, \emph{{Quintessential Starobinsky inflation and swampland
  criteria}},  \href{https://arxiv.org/abs/1810.06445}{{\ttfamily 1810.06445}}.

\bibitem{Fukuda:2018haz}
H.~Fukuda, R.~Saito, S.~Shirai and M.~Yamazaki, \emph{{Phenomenological
  Consequences of the Refined Swampland Conjecture}},
  \href{https://arxiv.org/abs/1810.06532}{{\ttfamily 1810.06532}}.

\bibitem{Cecotti:2015wqa}
S.~Cecotti, \emph{{Supersymmetric Field Theories}}. Cambridge University Press,
  2015.

\bibitem{Palti:2015xra}
E.~Palti, \emph{{On Natural Inflation and Moduli Stabilisation in String
  Theory}}, \href{https://doi.org/10.1007/JHEP10(2015)188}{\emph{JHEP}
  {\bfseries 10} (2015) 188}
  [\href{https://arxiv.org/abs/1508.00009}{{\ttfamily 1508.00009}}].

\bibitem{Baume:2016psm}
F.~Baume and E.~Palti, \emph{{Backreacted Axion Field Ranges in String
  Theory}}, \href{https://doi.org/10.1007/JHEP08(2016)043}{\emph{JHEP}
  {\bfseries 08} (2016) 043}
  [\href{https://arxiv.org/abs/1602.06517}{{\ttfamily 1602.06517}}].

\bibitem{Klaewer:2016kiy}
D.~Klaewer and E.~Palti, \emph{{Super-Planckian Spatial Field Variations and
  Quantum Gravity}}, \href{https://doi.org/10.1007/JHEP01(2017)088}{\emph{JHEP}
  {\bfseries 01} (2017) 088}
  [\href{https://arxiv.org/abs/1610.00010}{{\ttfamily 1610.00010}}].

\bibitem{Valenzuela:2016yny}
I.~Valenzuela, \emph{{Backreaction Issues in Axion Monodromy and Minkowski
  4-forms}}, \href{https://doi.org/10.1007/JHEP06(2017)098}{\emph{JHEP}
  {\bfseries 06} (2017) 098}
  [\href{https://arxiv.org/abs/1611.00394}{{\ttfamily 1611.00394}}].

\bibitem{Blumenhagen:2017cxt}
R.~Blumenhagen, I.~Valenzuela and F.~Wolf, \emph{{The Swampland Conjecture and
  F-term Axion Monodromy Inflation}},
  \href{https://doi.org/10.1007/JHEP07(2017)145}{\emph{JHEP} {\bfseries 07}
  (2017) 145} [\href{https://arxiv.org/abs/1703.05776}{{\ttfamily
  1703.05776}}].

\bibitem{Palti:2017elp}
E.~Palti, \emph{{The Weak Gravity Conjecture and Scalar Fields}},
  \href{https://doi.org/10.1007/JHEP08(2017)034}{\emph{JHEP} {\bfseries 08}
  (2017) 034} [\href{https://arxiv.org/abs/1705.04328}{{\ttfamily
  1705.04328}}].

\bibitem{Lust:2017wrl}
D.~L{\"u}st and E.~Palti, \emph{{Scalar Fields, Hierarchical UV/IR Mixing and
  The Weak Gravity Conjecture}},
  \href{https://doi.org/10.1007/JHEP02(2018)040}{\emph{JHEP} {\bfseries 02}
  (2018) 040} [\href{https://arxiv.org/abs/1709.01790}{{\ttfamily
  1709.01790}}].

\bibitem{Hebecker:2017lxm}
A.~Hebecker, P.~Henkenjohann and L.~T. Witkowski, \emph{{Flat Monodromies and a
  Moduli Space Size Conjecture}},
  \href{https://doi.org/10.1007/JHEP12(2017)033}{\emph{JHEP} {\bfseries 12}
  (2017) 033} [\href{https://arxiv.org/abs/1708.06761}{{\ttfamily
  1708.06761}}].

\bibitem{Cicoli:2018tcq}
M.~Cicoli, D.~Ciupke, C.~Mayrhofer and P.~Shukla, \emph{{A Geometrical Upper
  Bound on the Inflaton Range}},
  \href{https://doi.org/10.1007/JHEP05(2018)001}{\emph{JHEP} {\bfseries 05}
  (2018) 001} [\href{https://arxiv.org/abs/1801.05434}{{\ttfamily
  1801.05434}}].

\bibitem{Grimm:2018ohb}
T.~W. Grimm, E.~Palti and I.~Valenzuela, \emph{{Infinite Distances in Field
  Space and Massless Towers of States}},
  \href{https://doi.org/10.1007/JHEP08(2018)143}{\emph{JHEP} {\bfseries 08}
  (2018) 143} [\href{https://arxiv.org/abs/1802.08264}{{\ttfamily
  1802.08264}}].

\bibitem{Heidenreich:2018kpg}
B.~Heidenreich, M.~Reece and T.~Rudelius, \emph{{Emergence of Weak Coupling at
  Large Distance in Quantum Gravity}},
  \href{https://doi.org/10.1103/PhysRevLett.121.051601}{\emph{Phys. Rev. Lett.}
  {\bfseries 121} (2018) 051601}
  [\href{https://arxiv.org/abs/1802.08698}{{\ttfamily 1802.08698}}].

\bibitem{Blumenhagen:2018nts}
R.~Blumenhagen, D.~Kl{\"a}wer, L.~Schlechter and F.~Wolf, \emph{{The Refined
  Swampland Distance Conjecture in Calabi-Yau Moduli Spaces}},
  \href{https://doi.org/10.1007/JHEP06(2018)052}{\emph{JHEP} {\bfseries 06}
  (2018) 052} [\href{https://arxiv.org/abs/1803.04989}{{\ttfamily
  1803.04989}}].

\bibitem{Lee:2018urn}
S.-J. Lee, W.~Lerche and T.~Weigand, \emph{{Tensionless Strings and the Weak
  Gravity Conjecture}},  \href{https://arxiv.org/abs/1808.05958}{{\ttfamily
  1808.05958}}.

\bibitem{Dine:1985he}
M.~Dine and N.~Seiberg, \emph{{Is the Superstring Weakly Coupled?}},
  \href{https://doi.org/10.1016/0370-2693(85)90927-X}{\emph{Phys. Lett.}
  {\bfseries 162B} (1985) 299}.

\bibitem{Kachru:2003aw}
S.~Kachru, R.~Kallosh, A.~D. Linde and S.~P. Trivedi, \emph{{De Sitter vacua in
  string theory}},
  \href{https://doi.org/10.1103/PhysRevD.68.046005}{\emph{Phys. Rev.}
  {\bfseries D68} (2003) 046005}
  [\href{https://arxiv.org/abs/hep-th/0301240}{{\ttfamily hep-th/0301240}}].

\bibitem{Balasubramanian:2005zx}
V.~Balasubramanian, P.~Berglund, J.~P. Conlon and F.~Quevedo,
  \emph{{Systematics of moduli stabilisation in Calabi-Yau flux
  compactifications}},
  \href{https://doi.org/10.1088/1126-6708/2005/03/007}{\emph{JHEP} {\bfseries
  03} (2005) 007} [\href{https://arxiv.org/abs/hep-th/0502058}{{\ttfamily
  hep-th/0502058}}].

\bibitem{Veneziano:2001ah}
G.~Veneziano, \emph{{Large N bounds on, and compositeness limit of, gauge and
  gravitational interactions}},
  \href{https://doi.org/10.1088/1126-6708/2002/06/051}{\emph{JHEP} {\bfseries
  06} (2002) 051} [\href{https://arxiv.org/abs/hep-th/0110129}{{\ttfamily
  hep-th/0110129}}].

\bibitem{ArkaniHamed:2005yv}
N.~Arkani-Hamed, S.~Dimopoulos and S.~Kachru, \emph{{Predictive landscapes and
  new physics at a TeV}},
  \href{https://arxiv.org/abs/hep-th/0501082}{{\ttfamily hep-th/0501082}}.

\bibitem{Dvali:2007hz}
G.~Dvali, \emph{{Black Holes and Large N Species Solution to the Hierarchy
  Problem}}, \href{https://doi.org/10.1002/prop.201000009}{\emph{Fortsch.
  Phys.} {\bfseries 58} (2010) 528}
  [\href{https://arxiv.org/abs/0706.2050}{{\ttfamily 0706.2050}}].

\bibitem{Dvali:2007wp}
G.~Dvali and M.~Redi, \emph{{Black Hole Bound on the Number of Species and
  Quantum Gravity at LHC}},
  \href{https://doi.org/10.1103/PhysRevD.77.045027}{\emph{Phys. Rev.}
  {\bfseries D77} (2008) 045027}
  [\href{https://arxiv.org/abs/0710.4344}{{\ttfamily 0710.4344}}].

\bibitem{talk-madrid}
M.~Reece, \emph{{Large Field Ranges, UV Cutoffs, and Photon Masses}},  talk at
  \emph{Vistas over the Swampland}, 2018,
  \href{https://workshops.ift.uam-csic.es/swampland}{https://workshops.ift.uam-csic.es/swampland}.

\bibitem{Ashoorioon:2011aa}
A.~Ashoorioon, U.~Danielsson and M.~M. Sheikh-Jabbari, \emph{{1/N resolution to
  inflationary $\eta$-problem}},
  \href{https://doi.org/10.1016/j.physletb.2012.06.034}{\emph{Phys. Lett.}
  {\bfseries B713} (2012) 353}
  [\href{https://arxiv.org/abs/1112.2272}{{\ttfamily 1112.2272}}].

\bibitem{Conlon:2012tz}
J.~P. Conlon, \emph{{Quantum Gravity Constraints on Inflation}},
  \href{https://doi.org/10.1088/1475-7516/2012/09/019}{\emph{JCAP} {\bfseries
  1209} (2012) 019} [\href{https://arxiv.org/abs/1203.5476}{{\ttfamily
  1203.5476}}].

\bibitem{Kaloper:2015jcz}
N.~Kaloper, M.~Kleban, A.~Lawrence and M.~S. Sloth, \emph{{Large Field
  Inflation and Gravitational Entropy}},
  \href{https://doi.org/10.1103/PhysRevD.93.043510}{\emph{Phys. Rev.}
  {\bfseries D93} (2016) 043510}
  [\href{https://arxiv.org/abs/1511.05119}{{\ttfamily 1511.05119}}].

\bibitem{Andriolo:2018lvp}
S.~Andriolo, D.~Junghans, T.~Noumi and G.~Shiu, \emph{{A Tower Weak Gravity
  Conjecture from Infrared Consistency}},
  \href{https://doi.org/10.1002/prop.201800020}{\emph{Fortsch. Phys.}
  {\bfseries 66} (2018) 1800020}
  [\href{https://arxiv.org/abs/1802.04287}{{\ttfamily 1802.04287}}].

\bibitem{Strominger:1996it}
A.~Strominger, S.-T. Yau and E.~Zaslow, \emph{{Mirror symmetry is T duality}},
  \href{https://doi.org/10.1016/0550-3213(96)00434-8}{\emph{Nucl. Phys.}
  {\bfseries B479} (1996) 243}
  [\href{https://arxiv.org/abs/hep-th/9606040}{{\ttfamily hep-th/9606040}}].

\bibitem{Silverstein:2008sg}
E.~Silverstein and A.~Westphal, \emph{{Monodromy in the CMB: Gravity Waves and
  String Inflation}},
  \href{https://doi.org/10.1103/PhysRevD.78.106003}{\emph{Phys. Rev.}
  {\bfseries D78} (2008) 106003}
  [\href{https://arxiv.org/abs/0803.3085}{{\ttfamily 0803.3085}}].

\bibitem{McAllister:2008hb}
L.~McAllister, E.~Silverstein and A.~Westphal, \emph{{Gravity Waves and Linear
  Inflation from Axion Monodromy}},
  \href{https://doi.org/10.1103/PhysRevD.82.046003}{\emph{Phys. Rev.}
  {\bfseries D82} (2010) 046003}
  [\href{https://arxiv.org/abs/0808.0706}{{\ttfamily 0808.0706}}].

\bibitem{Marchesano:2014mla}
F.~Marchesano, G.~Shiu and A.~M. Uranga, \emph{{F-term Axion Monodromy
  Inflation}}, \href{https://doi.org/10.1007/JHEP09(2014)184}{\emph{JHEP}
  {\bfseries 09} (2014) 184} [\href{https://arxiv.org/abs/1404.3040}{{\ttfamily
  1404.3040}}].

\bibitem{Blumenhagen:2014gta}
R.~Blumenhagen and E.~Plauschinn, \emph{{Towards Universal Axion Inflation and
  Reheating in String Theory}},
  \href{https://doi.org/10.1016/j.physletb.2014.08.007}{\emph{Phys. Lett.}
  {\bfseries B736} (2014) 482}
  [\href{https://arxiv.org/abs/1404.3542}{{\ttfamily 1404.3542}}].

\bibitem{Hebecker:2014eua}
A.~Hebecker, S.~C. Kraus and L.~T. Witkowski, \emph{{D7-Brane Chaotic
  Inflation}},
  \href{https://doi.org/10.1016/j.physletb.2014.08.028}{\emph{Phys. Lett.}
  {\bfseries B737} (2014) 16}
  [\href{https://arxiv.org/abs/1404.3711}{{\ttfamily 1404.3711}}].

\bibitem{Blumenhagen:2018hsh}
R.~Blumenhagen, \emph{{Large Field Inflation/Quintessence and the Refined
  Swampland Distance Conjecture}},
  \href{https://doi.org/10.22323/1.318.0175}{\emph{PoS} {\bfseries CORFU2017}
  (2018) 175} [\href{https://arxiv.org/abs/1804.10504}{{\ttfamily
  1804.10504}}].

\end{thebibliography}\endgroup

\pdfoutput=1
\documentclass[12pt]{article}
\usepackage[margin=.99in]{geometry}

\usepackage{amsmath,dsfont}
\usepackage{amssymb}
\usepackage{graphicx,color}
\usepackage{slashed,cite,xcolor}
\usepackage{hyperref}

\usepackage{ifpdf}

\newcommand{\be}{\begin{equation}}
\newcommand{\ee}{\end{equation}}
\newcommand{\ba}{\begin{eqnarray}}
\newcommand{\ea}{\end{eqnarray}}

\title{{\bf \bigskip \Large{The asymptotic dS Swampland Conjecture -\\ a simplified derivation and a potential loophole}}\bigskip}

\author{Arthur Hebecker$^a$ and Timm Wrase$^b$\bigskip\bigskip\\
\small $^a$Institute for Theoretical Physics, University of Heidelberg, \\ \small Philosophenweg 19, D-69120 Heidelberg, Germany\bigskip\\
\small $^b$Institute for Theoretical Physics, TU Wien,\\ \small Wiedner Hauptstrasse 8-10/136, A-1040 Vienna, Austria\bigskip
}

\date{}

\begin{document}

\maketitle

\abstract{
\noindent Recently, arguments for a refined de Sitter conjecture were put forward in \cite{Ooguri:2018wrx}. Using the large distance conjecture of \cite{Ooguri:2006in}, the authors provide evidence for this dS conjecture in asymptotic regimes of field space, where the parametric control of string theory becomes arbitrarily good. Their main tool is Bousso's covariant entropy bound \cite{Bousso:1999xy}. Here, we describe a simpler way to reach a similar conclusion. The argument is based on the fact that the cutoff of an effective theory with gravity decreases as the number of species grows. We then discuss a loophole in this argument and its possible counterpart in the assumptions underlying the entropy-based derivation. The idea is to consider potentials which, while they remain below an exponentially falling bound, have small oscillations leading locally to relatively flat regions or even to an infinite series of dS minima.
}

\thispagestyle{empty}


\newpage


\section{Introduction}
Following the proposal of the so called dS swampland conjecture~\cite{Obied:2018sgi} (see \cite{Brennan:2017rbf, Danielsson:2018ztv} for related reviews), a constructive debate concerning dS models in string theory has been revived. Moreover, the related phenomenology of dark energy and inflation has received much attention, see e.g.~\cite{Agrawal:2018own, Andriot:2018wzk, Dvali:2018fqu, Achucarro:2018vey, Garg:2018reu, Lehners:2018vgi, Kehagias:2018uem, Dias:2018ngv, Denef:2018etk, Colgain:2018wgk, Paban:2018ole, Roupec:2018mbn, Andriot:2018ept, Matsui:2018bsy, Ben-Dayan:2018mhe, Heisenberg:2018yae, Damian:2018tlf, Conlon:2018eyr, Kinney:2018nny, Dasgupta:2018rtp, Cicoli:2018kdo, Kachru:2018aqn, Akrami:2018ylq, Murayama:2018lie, Marsh:2018kub, Choi:2018rze, DAmico:2018mnx, Han:2018yrk, Moritz:2018ani, Halverson:2018cio, Anguelova:2018vyr, Lin:2018kjm, Hamaguchi:2018vtv, Ashoorioon:2018sqb, Wang:2018kly, Fukuda:2018haz}. Clearly, the realization of dS minima and of inflation in string theory is a difficult and important subject and such a revived attention is very desirable.
An important point in our context is the suggestion of potential counter-examples to the original dS swampland conjecture in \cite{Denef:2018etk, Conlon:2018eyr, Cicoli:2018kdo, Murayama:2018lie, Choi:2018rze, Hamaguchi:2018vtv}. The refined version of the dS swampland conjecture advertised in \cite{Ooguri:2018wrx}, which is very close to that of~\cite{Garg:2018reu}, does not have these issues (see \cite{Dvali:2018fqu, Andriot:2018wzk} for earlier suggestions for refinement). Very roughly, the statement of the refined conjecture is that $|V'|/V>{\cal O}(1)$ whenever the potential is sufficiently strongly convex. The asymptotic form claims only that this holds as one moves to infinity in field space.

In \cite{Ooguri:2018wrx} the authors also point out an interesting connection between their refined dS swampland conjecture and the swampland distance conjecture \cite{Ooguri:2006in, Cecotti:2015wqa, Palti:2015xra, Baume:2016psm, Klaewer:2016kiy, Valenzuela:2016yny, Blumenhagen:2017cxt, Palti:2017elp, Lust:2017wrl, Hebecker:2017lxm, Cicoli:2018tcq, Grimm:2018ohb, Heidenreich:2018kpg, Blumenhagen:2018nts, Lee:2018urn}. They do so by using  Bousso's covariant entropy bound \cite{Bousso:1999xy} and thereby provide support for their refined dS swampland conjecture in the asymptotic regime, which is parametrically controlled in a strict, mathematical sense. This is a generalization of the old Dine-Seiberg argument \cite{Dine:1985he}.
This result of \cite{Ooguri:2018wrx} is also consistent with the well-known fact (recalled e.g. in the first paragraph of \cite{Denef:2018etk}) that dS vacua must live in a regime where corrections cannot be made arbitrarily small since a single leading (non-constant) term in the scalar potential always gives rise to a runaway. This is also consistent with existing constructions of dS vacua like the KKLT \cite{Kachru:2003aw} and LVS scenario \cite{Balasubramanian:2005zx}, where several terms compete with each other to give rise to dS solutions. 

While the authors of \cite{Ooguri:2018wrx} also consider the possible application of their refined dS swampland conjecture to the entire moduli space, we will focus here only on the asymptotic region, i.e., on the limit where we approach the boundary of field space. We discuss a simple argument, based on the species bound~\cite{Veneziano:2001ah, ArkaniHamed:2005yv, Dvali:2007hz, Dvali:2007wp}, why one expects that the potential decays exponentially in any such direction.\footnote{We have been informed that the authors of~\cite{Ooguri:2018wrx} were aware of the restrictions imposed by the species bound in their context. The connection between the species bound and the Dine-Seiberg argument has also been previously pointed out by Matt Reece~\cite{talk-madrid}. For earlier discussions of the species and entropy bounds in the context of quasi-de-Sitter solutions see e.g.~\cite{Ashoorioon:2011aa,Conlon:2012tz, Kaloper:2015jcz}.}
While this asymptotic behavior seems generically consistent with the refined dS swampland conjecture, we point out a simple loophole based on an exponentially bounded potential that nevertheless has dS vacua. We then analyze how this loophole might show up in the assumptions of the entropy-based derivation. It would be interesting to check whether such a potential can or cannot arise in string theory.

\section{A simplified derivation}
We now use the distance conjecture \cite{Ooguri:2006in, Baume:2016psm, Klaewer:2016kiy} and the reduced cutoff in the presence of a large number of species \cite{Veneziano:2001ah, ArkaniHamed:2005yv, Dvali:2007hz, Dvali:2007wp} to show that the scalar potential should decay exponentially when we approach the boundary in field space.\footnote{For other applications of towers of light states in related contexts see, e.g.,~\cite{Heidenreich:2018kpg, Andriolo:2018lvp}.} Concretely, we argue that for canonical $\phi$ the scalar potential has to satisfy $V < e^{-a \phi}$ for $\phi\rightarrow \infty$, where $a$ is an order one parameter.

Let us focus on a single tower of light states and assume equal spacing, as in the case of KK or string winding modes. In other words, we assume that the masses are given by $n \cdot m$, where $n\in \mathbb{N}$ and $m\equiv\, M_P  e^{-b \phi}$ with some number $b$ that is expected to be order one according to the distance conjecture. The number of particles with a mass below the cutoff is given by $N \approx \Lambda/m$. Next, similarly to \cite{Grimm:2018ohb}, we combine this with the well-known result \cite{Dvali:2007wp} that the cutoff of an effective gravitational theory is reduced in the presence of a large number of species, $\Lambda \approx M_P /\sqrt{N}$. Eliminating $N$ from these two relations, we find the highest value of $\Lambda$ for which the theory with this tower of states remains perturbatively controlled:
\be
\Lambda \approx (M_P^2 m)^{\frac13} \approx M_P e^{-\frac{b\phi}{3}}\,.
\ee 
We therefore find that the cutoff $\Lambda$ decreases exponentially, with the order one pre-factor $b/3$. In the case of multiple towers or, more precisely, a higher-dimensional lattice of light states, the prefactor changes but the crucial exponential behavior will remain.

Now let us allow for a positive scalar potential $V(\phi)$. Such a potential will in general (with details depending on its steepness and the overall dynamics) introduce a curvature scale $\sim H$, with $H$ defined by $3H^2M_P^2=V$. To have perturbative control, we need an effective field theory regime at a scale above $H$, i.e., we need $\Lambda\gtrsim H$. In other words, 
\be
M_P^2e^{-2b\phi/3}\gtrsim V(\phi)/M_P^2\,.
\ee
which is our desired exponential bound on the potential.

This result differs in two points from the one obtained in \cite{Ooguri:2018wrx}. Firstly, it is only an upper bound while the argument in \cite{Ooguri:2018wrx} gives the actual behavior of the potential. Secondly, it relies only on the reduced cutoff and the distance conjecture. By contrast, the derivation in \cite{Ooguri:2018wrx} requires an apparent horizon. This constrains in particular the second derivative of the scalar potential, which hence motivated the refined dS swampland conjecture in \cite{Ooguri:2018wrx}.

\section{A potential loophole}
We have provided evidence, in an arguably simpler way than that of \cite{Ooguri:2018wrx}, for the exponential decay of the scalar potential. Let us now assume that indeed $V(\phi) < e^{-a \phi}$ as one approaches the boundary of field space, $\phi \rightarrow \infty$. One would like to further conclude that $|V'|/V\gtrsim a$ since the potential decays with an exponent $-a$ or faster. While we do expect this behavior in string theory, it is unfortunately not so easy to derive this last result in a strict sense. Indeed, consider the following simple counter-example:

If we multiply the exponentially decaying potential by an oscillating function, e.g.
\be
V(\phi)=e^{-a\phi}[A+B\cos(\alpha\phi)]\,,\label{pot}
\ee
then such a potential can have metastable dS vacua.\footnote{
See \cite{Palti:2017elp} for a different use of the idea of oscillations to establish a loophole in a Swampland-type argument.} Moreover, even imposing monotonicity as an extra ad hoc constraint, the desired conclusion cannot be reached. Given appropriately chosen values of the parameters in (\ref{pot}), one can ensure that $V$ falls monotonically but has infinitely many, arbitrarily flat regions. Such oscillations are compatible with the species-bound argument above.

They do not provide a loophole in the logic of \cite{Ooguri:2018wrx} but may rather correspond to some of the assumptions in \cite{Ooguri:2018wrx} being too strong. Specifically, the logic presented there postulates a monotonic behavior of the prefactor $n(\phi)$ in the exponential growth of species, $N(\phi)=n(\phi)\exp(c\phi)$, below a certain cutoff (which must also depend on $\phi$ and the definition of which is not a priori obvious). Then, this number $N$ of species has to saturate the dS entropy precisely.
\footnote{Let us note, independently of our main logic, that in controlled string models this would be very special. Indeed, the known asymptotic directions are large volume $(t\to \infty)$, large complex structure $(z\to \infty)$, and small $g_s$ ($g_s\to 0$). In the last case, the light tower are string excitations and stabilizing the compactification space above that scale is hard. By the Strominger-Yau-Zaslov conjecture \cite{Strominger:1996it}, going to large $z$ at fixed $t$ means going to large volume of the mirror dual. So we can focus on the latter case. But, at large volume, the light tower starts at the KK scale. Now (setting $M_P=1$) one may use the free-particle entropy relation $S\sim N^{1/4}R^{3/2}$ \cite{Ooguri:2018wrx}  for $N$ species from above $m_{KK}$. Together with $S\sim R^2$ and $R\sim 1/H$, this gives $H\sim 1/R\sim 1/\sqrt{N}$. But $1/\sqrt{N}$ is the species cutoff, implying $1/\sqrt{N}\gg m_{KK}$ and hence $H\gg m_{\rm KK}$. This differs from conventional settings where $H\ll m_{\rm Moduli}\ll m_{\rm KK}$ and it would be interesting to see how such an unusual regime can arise.}
Thus, to close the oscillations-loophole using the entropy argument, one would need a precise definition of the relevant number of species (via the cutoff or otherwise), and furthermore demonstrations that this number grows in a precisely exponential way (as above) {\it and} saturates the dS entropy precisely.  By precisley we here mean without any `wiggle room' of ${\cal O}(1)$ factors. We believe that more work is needed to understand these points and to exclude oscillations on general grounds.

Until then, it remains an interesting question whether such a scalar potential (or a variant thereof) can arise in string theory. It does not seem likely that saxion directions would receive such periodic corrections. However, one can break the shift symmetry along axionic directions~\cite{Silverstein:2008sg, McAllister:2008hb}, for example by a flux-induced $F$-term potential~\cite{Marchesano:2014mla, Blumenhagen:2014gta, Hebecker:2014eua}. The desired outcome of this was an axion-monodromy-type inflationary potential. At least in some cases, the appearance of the latter may be prevented by the backreaction of the axion on the saxion~\cite{Palti:2015xra, Baume:2016psm, Klaewer:2016kiy} (as reviewed in Sect.~5 of \cite{Blumenhagen:2018hsh}). Either way one always expects, based on the argument of the previous section, to find an asymptotically decaying potential that in this case could have oscillations. We leave it to the future to construct or rule out models with infinite directions in field space that involve the axion and saxion and give rise to such asymptotic oscillations.

\section{Conclusion}
We have discussed a simple argument for an exponential bound on scalar potentials at the boundary of moduli space: $V(\phi) < e^{-a \phi}$ for $\phi \rightarrow \infty$ and with $a\sim{\cal O}(1)$. The argument combines the appearance of many light fields according to the distance conjecture \cite{Ooguri:2006in} with the species bound \cite{Veneziano:2001ah, ArkaniHamed:2005yv, Dvali:2007hz, Dvali:2007wp}. The latter then limits the UV cutoff for $\phi\to \infty$. While this argument is rather straightforward and has probably been known to many of the experts, we find its discussion worthwhile for two reasons.

Firstly, it complements the derivation in~\cite{Ooguri:2018wrx}, where the authors argue for the actual behavior of the scalar potential: $V \sim e^{-b\phi}$. This result requires assumptions about or the knowledge of several ${\cal O}(1)$ factors in the interplay between cutoff, species number and de Sitter entropy. From the perspective of our simpler argument these assumptions correspond to the loophole of an exponentially bounded but oscillating potentials. In this sense, our note highlights a particularly non-trivial and possibly not yet finalized aspect of the entropy-based derivation of $V \sim e^{-b\phi}$.

Secondly, by suggesting a concrete way in which de Sitter or slow-roll regions for $\phi\to \infty$ can arise (e.g.~via backreacted axion-monodromy potentials), we open up the possibility of trying to challenge the asymptotic de Sitter conjecture in a constructive way. It may be an interesting exercise to attempt the realization of such scalar potentials in string theory. Alternatively, one can also prove their non-existence by establishing the assumptions on which the derivation of \cite{Ooguri:2018wrx} is based. Both of avenues provide important and feasible research projects for the future.

\vspace*{.2cm}
\noindent
{\bf Acknowledgments:} 
We would like to thank Philipp Henkenjohann, Eran Palti and Pablo Soler for useful discussions. TW is supported by an FWF grant with the number P 30265. 

\bibliographystyle{JHEP}
\bibliography{refs}

\end{document}